OpenStudy (anonymous):

i have a weird question maybe someone can help.... q^2+1...how do u factor?

OpenStudy (anonymous):

you cannot with real numbers

OpenStudy (anonymous):

That's true..

OpenStudy (anonymous):

so does that mean its prime?

OpenStudy (anonymous):

\[(q-1)(q-1)\]

OpenStudy (anonymous):

that is not correct

OpenStudy (anonymous):

No, if you are to solve for q, you have \[\sqrt{-1}=\sqrt{i}\] This requires that you are familiar with complex numbers.

OpenStudy (anonymous):

but then the middle would be -2q

OpenStudy (anonymous):

\[(q+1)(q+1)\]

OpenStudy (anonymous):

then be +2

OpenStudy (anonymous):

i mean +2q

myininaya (myininaya):

does (q-i)(q+i) work

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no cause then it would be -1 myininaya

myininaya (myininaya):

q^2+iq-iq-i^2 q^2+0-(-1) q^2+1 yes it works ! :)

OpenStudy (anonymous):

but its q^2+1

OpenStudy (anonymous):

listen

OpenStudy (anonymous):

i=sqrt(-1)

myininaya (myininaya):

q^2+1=(q-i)(q+i)

OpenStudy (anonymous):

that is called a complex number

OpenStudy (anonymous):

has a lot of interesting bits

OpenStudy (anonymous):

-1*+1 is still -1

OpenStudy (anonymous):

nooooooooooooooooooooooooooooooooooooooooo

myininaya (myininaya):

(q-i)(q+i)=q^2-qi+qi-i^2 agreed?

OpenStudy (anonymous):

-i*i=-i^2=-(-1)=1

OpenStudy (anonymous):

its not i its 1(one)

myininaya (myininaya):

q^2+0-(-1) q^2+1

myininaya (myininaya):

i^2=-1

OpenStudy (anonymous):

\[(q+1)(q+1) = q ^{2}+1^{2}\]

myininaya (myininaya):

lol this is correct to say q^2+1=(q-i)(q+i)

OpenStudy (anonymous):

NOOOOO it is not

OpenStudy (anonymous):

either im super stupid or im just not comprehending how that works

myininaya (myininaya):

shark thats wrong (q+1)(q+1)=q^2+2q+1

OpenStudy (anonymous):

sharkattack you still get q^2 2q+1 as the answer

OpenStudy (anonymous):

As I said: you cannot with real numbers! (you need to introduce complex numbers with i to solve it)

OpenStudy (anonymous):

so andras does that mean its prime?

OpenStudy (anonymous):

\[(X+Y)(X+Y) = X ^{2} + Y ^{2}\]

OpenStudy (anonymous):

no it does not

myininaya (myininaya):

ok Bmartin i^2=-1 (q-i)(q+i)=q^2+qi-qi-i^2 =q^2-(-1) =q^2+1

OpenStudy (anonymous):

for example 3^2+1=10

myininaya (myininaya):

bmartin do you see the negative in front of the i^2?

OpenStudy (anonymous):

not prime

OpenStudy (anonymous):

so -1*+1 =+1?

OpenStudy (anonymous):

i dont believe that

OpenStudy (anonymous):

sorry but I'm right

myininaya (myininaya):

-(-1)=1

myininaya (myininaya):

i^2=-1 not 1

OpenStudy (anonymous):

myiniaya you are just confusing them with complex numbers

myininaya (myininaya):

you can factor over complex numbers

OpenStudy (anonymous):

Sharkattack is right!!! Lets vote him for president!!!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

so the answer is (q+1)(q-1)? i just really dont see how that works....

myininaya (myininaya):

no! it is (q-i)(q+i)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

no there is no answer

OpenStudy (anonymous):

why does it make a difference where the signs go?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

with normal (real) numbers there is no answer

OpenStudy (anonymous):

oooooooooooooh im soooooo confused

OpenStudy (anonymous):

I can imagine, what is not clear?

OpenStudy (anonymous):

\[(X+Y)(X+Y) =X ^{2} +Y ^{2}\]

myininaya (myininaya):

lol (q-i)(q+i)=q^2-qi+qi-i^2=q^2-(-1)=q^2+1 how is not one seeing this except me i dont want to be on a debate team accept my answer or not you can either say not factorable over the real numbers or you can say the answer is (q-i)(q+i)

OpenStudy (anonymous):

shark it is not: it is x^2+2xy+y^2

OpenStudy (anonymous):

\[(q+1)(q+1) =q ^{2} +1^{2} =q ^{2} +1\]

OpenStudy (anonymous):

i just keep seeing that answer as q^2-1 myininaya

OpenStudy (anonymous):

myininaya you are right but if someone never learnt about complex numbers than you cannot explain them in 2 min

myininaya (myininaya):

i^2=-1 -i^2=1

OpenStudy (anonymous):

they have the same base

myininaya (myininaya):

do you know the bmartin that -i^2=1?

OpenStudy (anonymous):

sharck when you expand a bracket you have to multiply every term with every term so (x+1)(x+1)=x^2+x+x+1

OpenStudy (anonymous):

mmmmm....

myininaya (myininaya):

bmartin, do you know that -i^2=1?

OpenStudy (anonymous):

no

myininaya (myininaya):

do you about imaginary numberts>

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i'm right in my law

myininaya (myininaya):

ok the sum of two real squares is not factorable over the reals

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

ok

myininaya (myininaya):

if you knew of imaginary numbers then we could factor it

OpenStudy (anonymous):

factorize x^2 +1 X^2+2xy+Y^2

OpenStudy (anonymous):

hopefully i get to that chapter soon

myininaya (myininaya):

lol sorry andras i thought they were talking about imagary numbers so i just figured they knew lol

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

well thanks everyone!

OpenStudy (anonymous):

this took longer than I thaught

OpenStudy (anonymous):

but

myininaya (myininaya):

by the way sean the square root of -1 is i not the square root of i

OpenStudy (anonymous):

yeah, true :-)

OpenStudy (anonymous):

I'm right

OpenStudy (anonymous):

lol

myininaya (myininaya):

shark is so cute

OpenStudy (anonymous):

you are as always, go for politician you would do great

OpenStudy (anonymous):

so persistant...you shoud be a lawyer

myininaya (myininaya):

i felt as i was debating against someone such tough debater

OpenStudy (anonymous):

i bet he was on the debate team!

OpenStudy (anonymous):

can we convince you shark that you are wrong?

myininaya (myininaya):

what about if we 9^2+1 is this equal to 9^2+9(1)+1^2